# PIETRA: Physics-Informed Evidential Learning for Traversing Out-of-Distribution Terrain

### v1

@@ Abstract
Supplementary Video: https://youtu.be/OTnNZ96oJRk
@@ I Introduction
Recent advancements in perception and mobility have accelerated the deployment of autonomous robots in challenging real-world environments such as office spaces, construction sites, forests, deserts and Mars [1, 2, 3, 4, 5, 6], where both geometric and semantic comprehension of the terrain is crucial for reliable navigation. In these settings, self-supervised traversability learning has emerged as a powerful tool to train neural networks (NNs) to predict terrain models from navigation data without manual labeling [7, 8, 9, 10, 11], where the learned representations can work directly with model-based motion planners, providing better interpretability and flexibility compared to fully learned navigation policies. However, the lack of abundant and diverse training data limits the reliability of learned traversability models in novel environments, such as the situation considered in this work (see Fig. 1). This is a well-known issue in the learning literature that arises from out-of-distribution (OOD) inputs at test time due to the distribution shift between training and test data [12, 13].
@@
Many recent works mitigate the risk of encountering OOD scenarios by quantifying the epistemic uncertainty, which is the model uncertainty due to distribution shift [13]. For example, OOD terrain can be detected by learning a density estimator for the training data to identify test input with low density [14, 15], or a terrain auto-encoder for detecting poorly reconstructed terrain [9, 16]. While avoiding OOD terrain has been shown to improve mission success rate in our prior work [11], doing so can be too conservative. To this end, we propose Physics-Informed Evidential TRAversability (PIETRA), a self-supervised learning framework that seamlessly combines learning-based and physics-based traversability analysis methods such that the downstream planner relies on the learned model for in-distribution (ID) terrain and the physics-based model for OOD terrain. Improving upon our prior work [11], we exploit the mathematical structure of evidential learning to embed a physics-based prior that automatically gets invoked when epistemic uncertainty is high. Moreover, we propose an uncertainty-aware physics-informed loss function inspired by existing works to regularize the learned model to improve generalization [17, 18].
@@ II Related Work
The field of traversability analysis studies how to infer suitability of terrain for navigation (see survey [19]). Compared to hand-crafting planning costs based on terrain features, directly learning traversability models from data requires less manual labeling and results in a more accurate assessment of vehicle-terrain interaction. Based on navigation data, visual and/or geometric features of the visited terrain can be used to train a traversability predictor based on estimated traversability values [20, 9, 21, 8, 22, 23]. Building upon this basic idea, visual and visual-inertial representation learning [10, 24], self-training with pseudo-labels for unvisited terrain [25], temporal fusion of robot states and sensor measurements [7], and data augmentation via vision foundation models [26] can all improve learning accuracy. Alternatively, when a hand-crafted traversability model such as [27] is available, it can be used to provide supervisory signals for training NNs that are faster [3, 2]. One concern is that, because learning-based methods rely on likely limited real-world data, the learned models may not generalize to environments unseen during training. While our work also uses self-supervised learning to obtain a traversability model, we incorporate physics knowledge into the model to improve generalization to OOD environments.
OOD detection is well studied and closely related to uncertainty quantification (see surveys [12, 13]). At a high level, input features that are not well-represented in the training data can lead to high epistemic uncertainty, which can be estimated with techniques such as Bayesian dropout [28], model ensembles [29], and evidential methods [30]. Epistemic uncertainty can also be estimated via a terrain auto-encoder for detecting high reconstruction error [16], a density estimator fit to the training data distribution [15], or Gaussian Process regression [31, 32]. Our prior work [11] adopts the evidential method proposed by [33] to efficiently identify OOD terrain and shows that avoiding OOD terrain during planning can improve mission success rate. However, always avoiding OOD terrain can be too conservative. To address this limitation, this new work exploits the mathematical formulation of evidential learning to explicitly embed a physics prior that is activated when encountering OOD features. While informative priors have recently been combined with evidential learning in other problem settings, such as the use of a rule-based prior for trajectory prediction in autonomous driving [34], our work focuses on off-road navigation and additionally introduces a physics-informed loss to further improve OOD generalization.
@@ IV Physics-Informed Evidential Learning
In this section, we present the proposed method PIETRA shown in Fig. 3. At a high level, for each traversability parameter, PIETRA outputs a categorical distribution over discretized traversability values to capture aleatoric uncertainty (inherent, irreducible data uncertainty). Moreover, PIETRA uses a normalizing flow [43] to estimate densities of latent features as proxies for epistemic uncertainty (model uncertainty due to distribution shift). PIETRA improves upon our prior work EVORA [11] by having an explicit physics prior that is invoked when encountering OOD inputs (Sec. IV-A) and an uncertainty-aware physics-informed loss that implicitly injects physics knowledge to further improve model accuracy in OOD terrain (Sec. IV-B). Lastly, we design custom physics priors used in our experiments in Sec. IV-C and discuss the implementation details in Sec. IV-D.
@@ IV-B Uncertainty-Aware Physics-Informed Loss
where ${\text{cs}}:\mathbb{R}^{B}\rightarrow\mathbb{R}^{B}$ is the cumulative sum operator, and $\mathbf{p}$ and $\mathbf{y}$ are the predicted and target PMFs. To train an evidential NN, our prior work proposes the uncertainty-aware ${\text{EMD}^{2}}$ (${\text{UEMD}^{2}}$) loss, which is the expected ${\text{EMD}^{2}}$ under the predicted Dirichlet $q$ (see the closed form in [11, Thm. 1]):
@@ IV-D Implementation Details
The traversability predictor $p_{\bm{\phi}}$ and the normalizing flow $p_{\bm{\lambda}}$ are trained jointly using the proposed UPI loss (14) on an empirically collected dataset $\{(\mathbf{o},\bm{\psi})_{k}\}_{k=1}^{K}$ where $K>0$ and every estimated traversability parameter in $\bm{\psi}$ is converted to a target PMF $\mathbf{y}$ using one-hot encoding. In practice, we assume independence of the traversability parameters and learn their distributions separately. For simplicity, we use a multi-layer perceptron (MLP) as the shared encoder to process the flattened and concatenated semantic and elevation patches of the terrain. The shared encoder is followed by separate fully connected decoders with soft-max outputs for predictions of each traversability parameter. To reduce computation, we train a single flow network using the outputs from the shared encoder. As the accuracy of physics priors differs across traversability parameters, we use a standalone MLP to map the shared encoder’s output to scalars between 0 and 1 to downscale the predicted evidence in the posterior update (11) for each traversability parameter.
@@ V-A Simulation Setup
An overview of the simulation environment is shown in Fig. 5. We generate 30 distinct synthetic elevation maps based on the real-world rock testbed proposed in [46], which are split evenly for training, validation and testing. Every map has sides of $50$ m and we use an Ackermann-steering robot of dimension $3.4$ m by $1.8$ m to collect 20 navigation trials in each map. For simplicity, the robot executes sinusoidal steering and $2$ m/s speed commands, and the episode ends if the robot rolls over, gets stuck, or reaches the map boundaries.
@@ V-B Learning Benchmark
The overall, ID and OOD test errors are reported in Table I, where we also include the uniform prior and the proposed physics prior. The takeaway is that PIETRA achieves the best overall, ID and OOD performance. To gain more intuition, the test errors are binned based on terrain unevenness in Fig. 6. Notably, while PI outperforms Vanilla and performs similarly to PIETRA on ID features, the improvement thanks to the physics-informed loss degrades as features become more OOD. On the other hand, PIETRA and EVORA fall back to their corresponding priors (physics prior and uniform prior) due to decreasing predicted evidence (11).
@@ V-D Navigation Benchmark
We deploy the NNs from Sec. V-B trained with the best hyperparameters in the test environments. As discussed in [11], EVORA detects OOD terrain that is avoided during planning via auxiliary costs. We consider 10 test maps and 10 start-goal pairs that are 40 m apart as visualized in Fig. 5. The robot aims to reach the goal without exceeding $30^{\circ}$ of roll and pitch angles or getting immobilized. The robot has a maximum steering angle of $30^{\circ}$ and a maximum speed of $1$ m/s to ensure stable crawling. We further consider 3 risk tolerances $\alpha$ in $\{0.4,0.6,0.8\}$ for the planner and report the success rate and the time to goal of successful trials.
The navigation results are shown in Fig. 7, which shows that PIETRA achieves the best success rate and time to goal. In comparison, EVORA is too conservative by always avoiding OOD terrain. While PI outperforms Vanilla, it performs worse than the physics prior due to poor OOD generalization. Note that the test maps are challenging to navigate, so the best average success rate is low, but the results still clearly demonstrate the advantages of the proposed method.
Figure 7: Success rate and time to goal from the simulated navigation benchmark. The figure shows the mean and standard deviation values for each method. The proposed PIETRA performs the best because it seamlessly transitions between the learned and physics models on OOD terrain. EVORA is too conservative because it avoids OOD terrain. While PI outperforms Vanilla, it performs worse than the physics prior due to unreliable traversability predictions.
@@ VI-A Data Collection, Training and Deployment
The indoor 9.6 m by 8 m arena contains turf and fake bushes to mimic outdoor vegetation. In addition, the dirt semantic type includes the concrete floor and a skateboard ramp. A 0.33 m by 0.25 m RC car carries a RealSense D455 depth camera, and a computer with an Intel Core i7 CPU and Nvidia RTX 2060 GPU. The robot runs onboard traversability prediction, motion planning, and elevation mapping with 0.1 m resolution, but Vicon is used for estimating the pose and velocity of the robot. The robot identifies vegetation by extracting green image pixels instead of using a standalone NN classifier to conserve GPU resources.
@@ VI-B Navigation Results
The robot is tasked to navigate to a fixed goal without violating roll and pitch constraints of $30^{\circ}$ and $45^{\circ}$ respectively, while deciding whether to drive on vegetation, climb over the ramp, or stay on the concrete floor. We vary the elevation and density of bushes and repeat each method 20 times. Some snapshots of the start of the missions are visualized in Fig. 8 for scenarios with short and tall vegetation. The success rates and the time to goal of the successful trials are reported in Fig. 9, showing that PIETRA achieves the best performance. PIETRA violates the roll constraint once because the robot drives off the down-ramp prematurely. EVORA has 3 out-of-bound failures because the robot frequently takes wide turns to avoid OOD terrain and does not have time to stop within the boundaries. The physics prior leads to 2 out-of-bound failures because it misses the goal region slightly due to over-optimistic traction estimates. The physics prior also leads to 2 roll violations due to driving off the down-ramp prematurely.
@@ VII Conclusion and Future Work
One of the limitations of PIETRA is its large memory footprint due to yaw dependency and fine discretization, so improvements are needed to use PIETRA for higher-dimensional systems with more complex state dependencies. There are several directions that could be pursued to further improve the OOD generalization, such as data augmentation [18], differentiable physics simulators [39] and neuro-symbolic networks [40].

### v2

@@ I Introduction
Recent advancements in perception and mobility have accelerated the deployment of autonomous robots in challenging real-world environments such as office spaces, construction sites, forests, deserts and Mars [1, 2, 3, 4, 5, 6], where both geometric and semantic comprehension of the terrain is crucial for reliable navigation. In these settings, self-supervised traversability learning has emerged as a powerful tool to train neural networks (NNs) to predict terrain models from navigation data without manual labeling [7, 8, 9, 10], where the learned representations can work directly with model-based motion planners, providing better interpretability and flexibility compared to fully learned navigation policies. However, the lack of abundant and diverse training data limits the reliability of learned traversability models in novel environments. This is a well-known issue in the learning literature that arises from out-of-distribution (OOD) inputs at test time due to the distribution shift between training and test data [11, 12].
@@
Many recent works mitigate the risk of encountering OOD scenarios by quantifying the epistemic uncertainty, which is the model uncertainty due to distribution shift [12]. For example, OOD terrain can be detected by learning a density estimator for the training data to identify test input with low density [13, 14], or a terrain auto-encoder for detecting poorly reconstructed terrain [8, 15]. While avoiding OOD terrain has been shown to improve mission success rate in our prior work [10], doing so can be too conservative, such as the situation considered in this work (see Fig. 1). To this end, we propose Physics-Informed Evidential TRAversability (PIETRA), a self-supervised learning framework that seamlessly combines learning-based and physics-based traversability analysis methods such that the downstream planner relies on the learned model for in-distribution (ID) terrain and the physics-based model for OOD terrain. Improving upon our prior work [10], we exploit the mathematical structure of evidential learning to embed a physics-based prior that automatically gets invoked when epistemic uncertainty is high. Moreover, we propose an uncertainty-aware physics-informed loss function inspired by existing works to regularize the learned model to improve generalization [16, 17].
@@ II Related Work
The field of traversability analysis studies how to infer suitability of terrain for navigation (see survey [18]). Compared to hand-crafting planning costs based on terrain features, directly learning traversability models from data requires less manual labeling and results in a more accurate assessment of vehicle-terrain interaction. Based on navigation data, visual and/or geometric features of the visited terrain can be used to train a traversability predictor based on estimated traversability values [19, 8, 20, 21, 22]. Building upon this basic idea, visual and visual-inertial representation learning [9, 23], self-training with pseudo-labels for unvisited terrain [24], temporal fusion of robot states and sensor measurements [7], and data augmentation via vision foundation models [25] can all improve learning accuracy. Alternatively, when a hand-crafted traversability model such as [26] is available, it can be used to provide supervisory signals for training NNs that are faster [3, 2]. One concern is that, because learning-based methods rely on likely limited real-world data, the learned models may not generalize to environments unseen during training. While our work also uses self-supervised learning to obtain a traversability model, we incorporate physics knowledge into the model to improve generalization to OOD environments.
OOD detection is well studied and closely related to uncertainty quantification (see surveys [11, 12]). At a high level, input features that are not well-represented in the training data can lead to high epistemic uncertainty, which can be estimated with techniques such as Bayesian dropout [27], model ensembles [28], and evidential methods [29]. Epistemic uncertainty can also be estimated via a terrain auto-encoder for detecting high reconstruction error [15], a density estimator fit to the training data distribution [14], or Gaussian Process regression [30]. Our prior work [10] adopts the evidential method proposed by [31] to efficiently identify OOD terrain and shows that avoiding OOD terrain during planning can improve mission success rate. However, always avoiding OOD terrain can be too conservative. To address this limitation, this new work exploits the mathematical formulation of evidential learning to explicitly embed a physics prior that is activated when encountering OOD features. While informative priors have recently been combined with evidential learning in other problem settings, such as the use of a rule-based prior for trajectory prediction in autonomous driving [32], our work focuses on off-road navigation and additionally introduces a physics-informed loss to further improve OOD generalization.
@@ IV Physics-Informed Evidential Learning
In this section, we present the proposed method PIETRA shown in Fig. 3. At a high level, the traversability predictor $p_{\bm{\phi}}$ parameterized by $\bm{\phi}$ outputs a categorical distribution over discretized traversability values to capture aleatoric uncertainty (inherent, irreducible data uncertainty) for each traversability parameter. Moreover, the normalizing flow network [41] $p_{\bm{\lambda}}$ parameterized by $\bm{\lambda}$ estimates the densities of latent features as proxies for epistemic uncertainty (model uncertainty due to distribution shift). PIETRA improves upon our prior work EVORA [10] by having an explicit physics prior that is invoked when encountering OOD inputs (Sec. IV-A) and an uncertainty-aware physics-informed loss that implicitly injects physics knowledge to further improve model accuracy in OOD terrain (Sec. IV-B). Lastly, we design custom physics priors used in our experiments in Sec. IV-C and discuss the implementation details in Sec. IV-D.
@@ IV-B Uncertainty-Aware Physics-Informed Loss
where ${\text{cs}}:\mathbb{R}^{B}\rightarrow\mathbb{R}^{B}$ is the cumulative sum operator, and $\mathbf{p}$ and $\mathbf{y}$ are the predicted and target PMFs. To train an evidential NN, our prior work proposes the uncertainty-aware ${\text{EMD}^{2}}$ (${\text{UEMD}^{2}}$) loss, which is the expected ${\text{EMD}^{2}}$ under the predicted Dirichlet $q$ (see the closed form in [10, Thm. 1]):
@@ IV-D Implementation Details
The traversability predictor $p_{\bm{\phi}}$ and the normalizing flow $p_{\bm{\lambda}}$ are trained jointly using the proposed UPI loss (14) on an empirically collected dataset $\{(\mathbf{o},\bm{\psi})_{k}\}_{k=1}^{K}$ where $K>0$ and every estimated traversability parameter in $\bm{\psi}$ is converted to a target PMF $\mathbf{y}$ using one-hot encoding with $B=12$ discrete bins. Terrain features can be extracted from a semantic elevation map (e.g., [43]), while traversability values can be computed using the pose and velocity estimates obtained from onboard odometry, along with the commanded velocities.
For simplicity, we use a multi-layer perceptron (MLP) as the shared encoder to process the flattened and concatenated semantic and elevation patches of the terrain. The shared encoder is followed by separate MLP decoders with soft-max outputs for predictions of each traversability parameter. To reduce computation, we train a single flow network using the encoder’s outputs. As recommended by [31], we scale the constant certainty budget $N$ exponentially with the dimension of the encoder’s outputs for numerical purposes. As the accuracy of physics priors differs across traversability parameters, we use a standalone MLP to map the shared encoder’s output to scalars between 0 and 1 to downscale the predicted evidence in the posterior update (11) for each traversability parameter.
@@ V-A Simulation Setup
An overview of the simulation environment is shown in Fig. 5. We generate 30 distinct synthetic elevation maps based on the real-world rock testbed proposed in [45], which are split evenly for training, validation and testing. Every map has sides of $50$ m and we use an Ackermann-steering robot of dimension $3.4$ m by $1.8$ m to collect 20 navigation trials in each map. For simplicity, the robot executes sinusoidal steering and $2$ m/s speed commands, and the episode ends if the robot rolls over, gets stuck, or exits the map. Note that terrain features and traversability values are obtained directly from the simulator.
@@ V-B Learning Benchmark
The learning benchmark results in Table I show that PIETRA achieves the best overall, ID and OOD performance. To gain more intuition, the test errors are binned based on terrain unevenness in Fig. 6. Notably, while PI outperforms Vanilla and performs similarly to PIETRA on ID features, the improvement thanks to the physics-informed loss degrades as features become more OOD. On the other hand, PIETRA and EVORA fall back to their corresponding priors (physics prior and uniform prior) due to decreasing predicted evidence (11).
@@ V-D Navigation Benchmark
We deploy the NNs from Sec. V-B trained with the best hyperparameters in the test environments. As discussed in [10], EVORA detects OOD terrain that is avoided during planning via auxiliary costs. As an ablation of the physics prior, we also consider PIETRA with OOD avoidance. We consider 10 test maps and 10 start-goal pairs that are 40 m apart as visualized in Fig. 5. The robot aims to reach the goal without exceeding $30^{\circ}$ of roll and pitch angles or getting immobilized. The robot has a maximum steering angle of $30^{\circ}$ and a maximum speed of $1$ m/s to ensure stable crawling. Note that MPPI runs at 10 Hz on GPU with a 5 s planning horizon and 1024 rollouts. We consider 3 risk tolerances $\alpha$ in $\{0.4,0.6,0.8\}$ and report the success rate and the time to goal of successful trials.
The navigation results in Fig. 7 show that PIETRA achieves the best success rate and time to goal. In comparison, PIETRA with OOD avoidance is more conservative but still outperforms EVORA thanks to the proposed UPI loss (14). While PI outperforms Vanilla, it performs worse than the physics prior due to poor OOD generalization.
Figure 7: Success rate and time to goal from the simulated navigation benchmark. The figure shows the mean and standard deviation values for each method. The proposed PIETRA performs the best because it seamlessly transitions between the learned and physics models on OOD terrain. When PIETRA avoids OOD terrain, it becomes more conservative but still outperforms EVORA thanks to the physics-informed training loss. While PI outperforms Vanilla, it performs worse than the physics prior due to unreliable traversability predictions.
@@ VI-A Data Collection, Training and Deployment
The indoor 9.6 m by 8 m arena contains turf and fake bushes to mimic outdoor vegetation. The dirt semantic type includes the concrete floor and a skateboard ramp. A 0.33 m by 0.25 m RC car carries a RealSense D455 depth camera, and a computer with an Intel Core i7 CPU and Nvidia RTX 2060 GPU. The robot runs onboard traversability prediction, motion planning, and elevation mapping with 0.1 m resolution, but Vicon is used for estimating the pose and velocity of the robot. For simplicity, terrain elevation and semantics are estimated separately based on depth and color measurements. Each map cell contains a single semantic type initially set to dirt and is later updated to vegetation if the corresponding pixels in the image space are green.
@@ VI-B Navigation Results
The robot is tasked to navigate to a fixed goal without leaving the bounding box of the arena or violating the roll and pitch constraints of $30^{\circ}$ and $45^{\circ}$ respectively, while deciding whether to drive on vegetation, climb over the ramp, or stay on the concrete floor. We vary the elevation and density of bushes and repeat each method 20 times. Some snapshots of the start of the missions are visualized in Fig. 8 for scenarios with short and tall vegetation. The success rates and the time to goal of the successful trials are reported in Fig. 9, showing that PIETRA achieves the best performance. PIETRA violates the roll constraint once because the robot drives off the down-ramp prematurely. EVORA has 3 out-of-bound failures because the robot frequently takes wide turns to avoid OOD terrain, which makes the robot more likely to miss the goal and leave the bounding box. The physics prior leads to 2 out-of-bound failures because it misses the goal region slightly due to over-optimistic traction estimates. The physics prior also leads to 2 roll violations due to driving off the down-ramp prematurely.
@@ VII Conclusion and Future Work
One of the limitations of PIETRA is its large memory footprint due to discretization, so improvements are needed for higher-dimensional systems. Moreover, PIETRA relies on accurate terrain features and state estimation, but the risk due to uncertain sensing and estimation may need to be addressed for some real-world environments. Lastly, PIETRA can be further improved by leveraging data augmentation [17], differentiable physics simulators [37] and neuro-symbolic networks [38].